OpenStudy (charlotte123):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

jimthompson5910 (jim_thompson5910):

all you need to do is find the slope of the line through (2, 13) and (−4, −11)

jimthompson5910 (jim_thompson5910):

this is because parallel lines have equal slopes

OpenStudy (charlotte123):

@jim_thompson5910 this is a pretest question so I don't really know how to do that yet T_T can u help me please? :)

jimthompson5910 (jim_thompson5910):

well if it's a pretest and you don't know it, then no big deal really

jimthompson5910 (jim_thompson5910):

if you know it, then you can answer it and move onto a higher level (and not have to relearn it) if you don't know it, then this will tell the teacher what you need to learn

jimthompson5910 (jim_thompson5910):

so either way, it works in your favor

OpenStudy (charlotte123):

@jim_thompson5910 exactly :) but I wanna get it right ^-^ can u give me an example of how to do this or help me in someway at least? :)

jimthompson5910 (jim_thompson5910):

I know you want to get it right, but it should be you that's doing this on your own I'll give you the formula and little bit after that, but will stop there

jimthompson5910 (jim_thompson5910):

The slope formula is m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

what this tells us is that the slope is equal to the change in y (y2 - y1) over the change in x (x2 - x1

OpenStudy (charlotte123):

thanks jim ^-^

OpenStudy (anonymous):

use y=mx+b

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

you will know y and x cuzz it is given to you and you know m from jims formula and you will have to find b :D

jimthompson5910 (jim_thompson5910):

also (x1, y1) = (2,13), so we can say x1 = 2 and y1 = 13 (x2, y2) = (-4,-11), so we can say x2 = -4 and y2 = -11 I'll let you take over

OpenStudy (charlotte123):

thanks again jim ^-^ and thanks timo :D

jimthompson5910 (jim_thompson5910):

sure thing